OpenStudy (anonymous):

Give a unique example of each: 1. Part 1: An absolute value equation with two solutions 2. Part 2: An absolute value equation with one solution 3. Part 3: An absolute value equation with no solutions

OpenStudy (anonymous):

1. \[|x|=3\] meaning \[x=\ \pm 3\]

OpenStudy (anonymous):

\[|x|=-2\]

OpenStudy (anonymous):

for part 3

OpenStudy (anonymous):

\[|2x+5|=4x-8\] \[2x+5=4x-8\] \[13=2x\] \[13/2=x\] \[|2x+5|=4x-8\] \[2x+5=-(4x-8)\] \[2x+5=-4x+8\] \[6x=3\] \[x=\frac{1}{2}\]

OpenStudy (anonymous):

|x| = 3 (2 solutions -3 and +3) |x| = 0 (x can only be 0) |x| = -5 (An absolute value can never be negative)